A study is conducted to investigate the impact of depression on a patient’s ability to survive cardiac disease. They evaluated 477 participants for depression and followed for 4 years. There were 374 patients with no depression (ND) and 69 of them died. Out of 103 patients with minor or major depression (D), 28 died. Construct a 95% confidence interval for the difference in proportion of deaths between depressed patients with cardiac disease and non-depressed patients with cardiac disease. (Hint: Round sample proportions and SE to four decimal places.)
Answer #1

Here, , n1 = 103 , n2 = 477
p1cap = 0.2718 , p2cap = 0.7841


Standard Error, sigma(p1cap - p2cap),
SE = sqrt(p1cap * (1-p1cap)/n1 + p2cap * (1-p2cap)/n2)
SE = sqrt(0.2718 * (1-0.2718)/103 + 0.7841*(1-0.7841)/477)
SE = 0.0477

For 0.95 CI, z-value = 1.96
Confidence Interval,
CI = (p1cap - p2cap - z*SE, p1cap - p2cap + z*SE)
CI = (0.2718 - 0.7841 - 1.96*0.0477, 0.2718 - 0.7841 + 1.96*0.0477)
CI = (-0.6058 , -0.4188)